A brand of television has a lifetime that is normally distributed with a mean of 7 years and a standard deviation of 2.5 years. What is the probability that a randomly chosen TV will last more than 8 years? Note: Write your answers with two places after the decimal, rounding off as appropriate.

Answer by stanbon(75887)

A brand of television has a lifetime that is normally distributed with a mean of 7 years and a standard deviation of 2.5 years. What is the probability that a randomly chosen TV will last more than 8 years? Note: Write your answers with two places after the decimal, rounding off as appropriate.
---------------
z(8) = (8-7)/2.5 = 1/2.5 = 0.4
------
P(x > 8) = P(z > 0.4) = normalcdf(0.4,100) = 0.3446
-------------
Cheers,
Stan H.
--------------